# Perturbations of Einstein-Maxwell-phantom spacetime: Instabilities of charged Ellis wormholes and quasinormal modes of black holes

### v1

@@ Abstract
Phantom scalar fields, as a viable candidate for dark energy, have been instrumental in eliminating spacetime singularities and constructing wormholes and regular black holes. We investigate the Einstein-Maxwell-phantom (EMP) framework, in which the Ellis wormholes can be charged and regular black holes can be admitted. While the previous study has shown the stability of EMP wormholes under massless scalar field perturbations, we further perform a comprehensive linear analysis of the EMP spacetime through phantom scalar field perturbations and gravito-electromagnetic field perturbations. Our analyses of effective potentials and finite difference time profiles reveal the linear instability of EMP wormholes. In the black hole scenario, the quasinormal spectra of the EMP theory, where the matrix-valued direct integration method and the Prony method are used, recover those of general relativity when the scalar charge goes to zero. By introducing the concepts of generalized specific charge and mixing angle, we quantify how the relative contributions between the phantom scalar and the electromagnetic fields produce distinct modifications in the quasinormal spectra.
@@ 1 Introduction
with $w<-1/3$ required for cosmic acceleration. Within the widely accepted $\Lambda$CDM model, dark energy is realized as the cosmological constant $\Lambda$, for which $w=-1$ [6]. Alternative scenarios, such as quintessence models [7, 8] and K-essence models [9, 10], allow $w>-1$, whereas phantom dark energy $(w<-1)$ would lead to a future Rip [11, 12, 13, 14, 15]. Although the current observations show remarkable consistency with $\Lambda$CDM, they do not exclude a slightly phantom behavior: several independent probes suggest $w\sim-1.03$ [16, 17, 18], and a combined analysis yields $w=-1.013^{+0.038}_{-0.043}$ [19].
In Einstein’s general theory of relativity, the acceleration of the universe can be accounted for either by adding a negative cosmological constant $\Lambda$ on the left-hand side of the Einstein equation or by introducing exotic matter with negative pressure to the source term on the right-hand side [12]. The former prescription realizes Gliner’s $\mu$-vacuum [20] in a de Sitter geometry, which can prevent spacetime singularity arising in gravitational collapse. A spacetime singularity, where the matter density and spacetime curvature diverge, marks the breakdown of physics. According to the Penrose singularity theorem [21], a singularity is inevitable within the framework of general relativity and under reasonable energy conditions, but no naked singularity exists, i.e., a singularity is hidden by event horizons. If the latter prescription is adopted, source terms that violate certain energy conditions are indispensable in order to construct regular (nonsingular) black holes. Notable examples include magnetically charged solutions in nonlinear electrodynamics [22, 23, 24] and configurations with phantom scalar hairs [25, 26, 27, 28, 29] (see Ref. [30] for a brief review).
The existence of black holes has been empirically confirmed through the historical direct detection of gravitational waves from a binary black hole merger by Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo [31], with numerous subsequent events reinforcing this statement [32, 33, 34, 35]. Moreover, Fermi Gamma-ray Burst Monitor (Fermi-GBM) and International Gamma-Ray Astrophysics Laboratory (INTEGRAL) observed [37, 36] a short gamma-ray burst from the same event in 1.74 seconds after LIGO and Virgo detected gravitational waves from the binary neutron star merger — GW170817. Furthermore, the first images of the supermassive black hole at the center of M87 [38, 39, 40, 41, 42, 43] and Sagittarius A* in the Milky Way [44, 45, 46, 47, 48, 49], captured by Event Horizon Telescope (EHT), inaugurate an era of multi-messenger astronomy. These breakthroughs provide us with more possibilities to test fundamental physical problems, e.g., the singularity problem [50, 51] and the validity of the no-hair theorem [52, 53, 54].
@@
As with regular black holes, wormhole models also require exotic matter that violates energy conditions, specifically the weak energy condition [80, 81]. The first eponymous wormhole solution was developed by Ellis [82] in a theory where the gravity is coupled to a massless phantom scalar field. In general, a phantom field is characterized by a flapping sign in the kinetic term of its action [83]. More recently, a charged Ellis wormhole was proposed [84] by the way to extend a class of Einstein-Maxwell-dilaton (EMD) theory [85] to an Einstein-Maxwell-phantom (EMP) theory through a complex transformation. In this EMP spacetime, sufficient electric charge yields a wormhole geometry, while below a critical charge the solution becomes a regular black hole, which can be classified as Type I or Type II based on its horizon structure. Further study [78] indicates that, under massless scalar field perturbations, EMP wormholes remain stable and exhibit echo signals near the wormhole-black hole transition threshold. Note that the EMP spacetime is sourced by the phantom scalar field and the Maxwell field, and the latter can also be phantom-like in certain region. Such phantom-like behaviors imply negative kinetic energy of the source terms and can potentially induce spacetime instabilities. Investigating both phantom scalar and gravito-electromagnetic field perturbations is therefore essential for assessing the dynamical stability of the EMP spacetime and may offer fresh insights into gravity.
In this work, we focus on the axial (gravito-electromagnetic) field perturbation of the EMP spacetime. The organization of this paper is as follows. In Sec. 2, we briefly review the derivation and properties of the EMP spacetime. In Sec. 3, we study the linear perturbations of the EMP spacetime and derive the master equations of phantom scalar field perturbations and gravito-electromagnetic field perturbations. In Sec. 4, we investigate the instability of EMP wormholes. In Sec. 5, we compute the quasinormal frequencies (QNFs) of Type I black holes using the matrix-valued direct integration method and the Prony method. Finally, the conclusion is made in Sec. 6.
2 Einstein-Maxwell-phantom spacetime
@@
where $g$ is the metric determinant, $R$ is the Ricci scalar and the dilaton scalar field $\phi_{\mathrm{d}}$ is minimally coupled to gravity, but nonminimally coupled to the Maxwell field $F=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}A$ in the form $\mathrm{e}^{-2\phi_{\mathrm{d}}}$. Rescaling the dilaton $\phi=2\phi_{\mathrm{d}}$ yields the standard coupling $\mathrm{e}^{-\phi}$. Moreover, such a nonminimal coupling is further generalized [85] to $\left(\gamma_{1}\cosh{\phi}+\gamma_{2}\sinh{\phi}\right)^{-1}$, where $\gamma_{1}$ and $\gamma_{2}$ are coupling constants, and the generalized coupling function reduces to $\mathrm{e}^{-\phi}$ when $\gamma_{1}=\gamma_{2}=1$. Under the following complex transformation,
@@
$\square\phi=\frac{1}{4}\frac{\partial Z^{-1}}{\partial\phi}F^{2},$ (5)
@@
where the energy-momentum tensors of the Maxwell field $T_{\mu\nu}^{A}$ and the phantom scalar field $T_{\mu\nu}^{\phi}$ take the forms,
@@
one gives [84] the charged Ellis wormhole solution of the EMP theory,
@@
where the regularization parameter $q$ corresponds to the radius of the wormhole throat and can be interpreted as the scalar charge; $Q$ is an integration constant that contributes to both the mass and electric charge of the EMP spacetime. Note that the EMP spacetime is asymptotically flat and regular, featuring a wormhole throat (also called a bounce) at $r=0$. The mass $M$ and electric charge $Q_{\mathrm{e}}$ are
@@
The horizon distribution of EMP black holes admits two distinct configurations: Type I resembles that of Reissner-Nordström (RN) black holes, where the two horizons locate on one side of the bounce; Type II possesses the horizon distribution of a bounce geometry that smoothly interpolates between the black hole and the white hole regions. The EMP framework reduces to general relativity in the classical limit,
$\gamma_{1}=\frac{1}{4},\qquad\gamma_{2}=\frac{8qM}{Q^{2}},\qquad q\to 0^{+}.$ (15)
This correspondence arises from the fundamental principles of general relativity, e.g., the preservation of U(1) gauge symmetry, enforcement of the no-scalar-hair theorem, and satisfaction of the positive mass theorem [86], thereby establishing theoretical consistency of the EMP theory with the Einstein-Maxwell dynamics.
@@ 3.1 General set up
The charged Ellis wormhole solution of the EMP theory [84], as a special case of Eq. (19), has the following coefficients,
@@
[…]
@@
3.2 Phantom scalar field perturbation
At first, we deal with the phantom scalar field perturbation. By linearizing the equation of motion Eq. (5), one obtains the master equation of phantom scalar field perturbations,
$\square\delta\phi$ $=\frac{1}{4}\left(\frac{\partial^{2}Z^{-1}}{{\partial\phi}^{2}}F^{2}\delta\phi+\frac{\partial Z^{-1}}{\partial\phi}\delta F^{2}\right).$ (28)
with the effective potential
$V=h\left(\frac{rh^{\prime}}{\varrho^{2}}+\frac{q^{2}h}{\varrho^{4}}+\frac{l(l+1)}{\varrho^{2}}+\frac{Q^{2}}{8q^{2}\varrho^{2}}\left(2rZ^{\prime}-\varrho^{2}Z^{\prime\prime}\right)\right),$ (31)
@@ 3.3 Gravito-Electromagnetic field perturbation
In the gravito-electromagnetic field perturbation, the linearization of the Einstein equation Eq. (7) yields
@@
The left-hand side of the above equation can be computed via Cartan’s equations of structure applied to the 1-forms in Eq. (22). In particular, the axial components $\delta R_{12}$ and $\delta R_{13}$ take the forms,
@@
where $\mathcal{Q}(r,\theta,t)=\varrho^{2}\mathrm{e}^{2\nu}Q_{23}\sin^{3}\theta$ encodes gravitational field perturbations and $B(r,\theta)=\delta P_{01}\sin{\theta}$ ) represents electromagnetic field perturbations. Note that those terms involving $\delta\left(\partial_{a}\phi\partial_{b}\phi\right)$ vanish in Eqs. (33) and (34) by virtue of the axial symmetry of the background metric Eq. (19). Eliminating $\sigma$ in Eqs. (33) and (34) and assuming a time dependence $\mathrm{e}^{\mathrm{i}\omega t}$ in the perturbations, we find the first master equation for gravito-electromagnetic field perturbations,
@@
$).\displaystyle\left[Z^{-1}\mathrm{e}^{2\nu}\left(Z\varrho\mathrm{e}^{\nu}B\right)_{,2}\right]_{,2}+\frac{\mathrm{e}^{\nu}}{\varrho}\left(\frac{B_{,3}}{\sin{\theta}}\right)_{,3}\sin{\theta}+\left(\omega^{2}\varrho\mathrm{e}^{-\nu}-\frac{Q^{2}}{\varrho^{3}}Z\mathrm{e}^{\nu}\right)B=-Q\frac{\mathcal{Q}_{,3}}{\varrho^{4}\sin{\theta}}.$ (37)
@@
so that $H_{1}$ and $H_{2}$ represent electromagnetic and gravitational field perturbations, respectively, with subscripts denoting the spin of the fields. We further introduce $H_{0}=\Psi(r)$ and adopt the tortoise coordinate $r_{*}$ defined by
@@
$\mathbf{H}=\left(\begin{array}[]{c}H_{0}\\ H_{1}\\ H_{2}\end{array}\right),\qquad\mathbf{V}=\left(\begin{array}[]{ccc}V&0&0\\ 0&V_{11}&V_{12}\\ 0&V_{21}&V_{22}\end{array}\right),$ (53)
@@
In the Reissner-Nordström (RN) background, the effective potential matrix in the master equation can be diagonalized by a constant matrix, yielding decoupled perturbation equations. However, the coupling cannot be removed in this manner [^1] It has come to our attention that some studies [88, 89, 90] have inappropriately applied the RN decoupling procedure in their work. To avoid further misleading, we explicitly clarify here that the RN decoupling procedure is not valid in the EMP spacetime. in the EMP spacetime, so we shall apply numerical methods directly to the coupled system.
4 Instability of wormholes

### v2

# Perturbations of Einstein–Maxwell–phantom spacetime: Instabilities of charged Ellis–Bronnikov wormholes and quasinormal modes of black holes

@@ Abstract
Phantom scalar fields, as a viable candidate for dark energy, have been instrumental in eliminating spacetime singularities and constructing wormholes and regular black holes. We investigate the Einstein–Maxwell–phantom (EMP) framework, in which the Ellis–Bronnikov wormholes can be charged and regular black holes can be admitted. While the previous study has shown the stability of EMP wormholes under massless scalar field perturbations, we further perform a comprehensive linear analysis of the EMP spacetime through gravito-electromagnetic field perturbations in the axial sector and phantom scalar field perturbations under an approximate treatment in the polar sector. Our analyses of effective potentials and finite difference time profiles reveal the linear instability of EMP wormholes. In the black hole scenario, the quasinormal spectra of Type I black holes, where the matrix-valued direct integration method and the Prony method are used, recover those of general relativity (GR) when the scalar charge goes to zero. Finally, by introducing the concepts of generalized specific charge and mixing angle, we quantify how the relative contributions between the phantom scalar and the electromagnetic fields modify the quasinormal spectra, and we assess the prospects for detecting spectral deviations between the EMP theory and GR in gravitational wave observation.
@@ 1 Introduction
[…]
In Einstein’s general theory of relativity, the acceleration of the universe can be accounted for either by adding a positive cosmological constant $\Lambda$ on the left-hand side of the Einstein equation or by introducing exotic matter with negative pressure to the source term on the right-hand side [12]. The former prescription realizes Gliner’s $\mu$-vacuum [20] in a de Sitter geometry, which can prevent spacetime singularity arising in gravitational collapse. A spacetime singularity, where the matter density and spacetime curvature diverge, marks the breakdown of physics. According to the Penrose singularity theorem [21], a singularity is inevitable within the framework of general relativity and under reasonable energy conditions, but the cosmic censorship conjecture states [22] that such singularities are always hidden behind event horizons, i.e., no naked singularities exist. If the latter prescription is adopted, source terms that violate certain energy conditions are indispensable in order to construct regular (nonsingular) black holes. Notable examples include magnetically charged solutions in nonlinear electrodynamics [23, 24, 25] and configurations with phantom scalar hairs [26, 27, 28, 29, 30] (see Ref. [31] for a brief review).
The existence of black holes has been empirically confirmed through the historical direct detection of gravitational waves from a binary black hole merger by Laser Interferometer Gravitational-Wave Observatory (LIGO) and Virgo [32], with numerous subsequent events reinforcing this statement [33, 34, 35, 36, 37]. Moreover, Fermi Gamma-ray Burst Monitor (Fermi-GBM) and International Gamma-Ray Astrophysics Laboratory (INTEGRAL) observed [39, 38] a short gamma-ray burst from the same event in 1.74 seconds after LIGO and Virgo detected gravitational waves from the binary neutron star merger — GW170817, inaugurating an era of multi-messenger astronomy. Furthermore, the first images of the supermassive black hole at the center of M87 [40, 41, 42, 43, 44, 45] and Sagittarius A* in the Milky Way [46, 47, 48, 49, 50, 51], captured by Event Horizon Telescope (EHT), provide direct high-resolution evidence for black holes in galactic nuclei. These breakthroughs provide us with more possibilities to test fundamental physical problems, e.g., the singularity problem [52, 53] and the validity of the no-hair theorem [54, 55, 56].
@@
As with regular black holes, wormhole models also require exotic matter that violates energy conditions, specifically the null energy condition (NEC) [82, 83]. The first eponymous wormhole solution — now known as the Ellis–Bronnikov (EB) wormhole — was developed independently by Ellis [84] and Bronnikov [85] in a theory where the gravity is coupled to a massless phantom scalar field. In general, a phantom field is characterized by a flapping sign in the kinetic term of its action [86]. More recently, a charged EB wormhole was proposed [87] by the way to extend a class of Einstein–Maxwell–dilaton (EMD) theory [88] to an Einstein–Maxwell–phantom (EMP) theory through a complex transformation. In this EMP spacetime, sufficient electric charge yields a wormhole geometry, while below a critical charge the solution becomes a regular black hole, which can be classified as Type I or Type II based on its horizon structure. Further study [80] indicates that, under massless scalar field perturbations, EMP wormholes remain stable and exhibit echo signals near the wormhole–black hole transition threshold, whereas instabilities of other phantom wormholes have been reported in Refs. [89, 90, 91, 92, 93]. Note that the EMP spacetime is sourced by the phantom scalar field and the Maxwell field, and the latter can also be phantom-like in certain region. Such phantom-like behaviors imply negative kinetic energy of the source terms and can potentially induce spacetime instabilities. Investigating both phantom scalar and gravito-electromagnetic field perturbations is therefore essential for assessing the dynamical stability of the EMP spacetime and may offer fresh insights into gravity.
The organization of this paper is as follows. In Sec. 2, we briefly review the derivation and key properties of the EMP spacetime. In Sec. 3, we study the linear perturbations of the EMP spacetime and derive the master equations of gravito-electromagnetic field perturbations in the axial sector and phantom scalar field perturbations in the polar sector, where the latter is discussed within a simplified and decoupled framework. In Sec. 4, we reveal instability of EMP wormholes by exhibiting the singular behavior in the gravito-electromagnetic effective potential and growth of the phantom scalar field perturbation. In Sec. 5, we compute the quasinormal frequencies (QNFs) of Type I black holes using the matrix-valued direct integration method and the Prony method, and we assess the observability of deviations between the gravitational QNMs of Type I black holes and Reissner–Nordström (RN) black holes. Finally, the conclusion is made in Sec. 6.
2 Einstein–Maxwell–phantom spacetime
@@
where $g$ is the metric determinant, $R$ is the Ricci scalar and the dilaton scalar field $\phi_{\mathrm{d}}$ is minimally coupled to gravity, but nonminimally coupled to the Maxwell field $F=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}A$ in the form $\mathrm{e}^{-2\phi_{\mathrm{d}}}$. Rescaling the dilaton $\phi=2\phi_{\mathrm{d}}$ yields the standard coupling $\mathrm{e}^{-\phi}$. Moreover, such a nonminimal coupling is further generalized [88] to $\left(\gamma_{1}\cosh{\phi}+\gamma_{2}\sinh{\phi}\right)^{-1}$, where $\gamma_{1}$ and $\gamma_{2}$ are coupling constants, and the generalized coupling function reduces to $\mathrm{e}^{-\phi}$ when $\gamma_{1}=\gamma_{2}=1$. A general classification of phantom EMD black holes has been presented in Ref. [94], where various causal structures and horizon configurations have been analyzed. The EMP theory considered here can be regarded as a particular extension of those phantom EMD models through the complex transformation,
@@
$\square\phi=-\frac{1}{4}\frac{\partial Z^{-1}}{\partial\phi}F^{2},$ (5)
@@
where the energy–momentum tensors of the Maxwell field $T_{\mu\nu}^{A}$ and the phantom scalar field $T_{\mu\nu}^{\phi}$ take the forms,
@@
one gives [87] the charged EB wormhole solution of the EMP theory,
@@
where the regularization parameter $q$ corresponds to the radius of the wormhole throat $Q$ is an integration constant that contributes to both the mass and electric charge of the EMP spacetime. Note that the EMP spacetime is asymptotically flat and regular, featuring a wormhole throat (also called a bounce) at $r=0$. The mass $M$ and electric charge $Q_{\mathrm{e}}$ are
@@
The horizon distribution of EMP black holes admits two distinct configurations:
  • Type I black hole. An outer event horizon and an inner Cauchy horizon lie on the same side of the wormhole throat (bounce); consequently the exterior causal structure closely resembles that of RN black holes.
  • Type II black hole. Two horizons straddle the throat, so that the bounce smoothly interpolates between black hole and white hole regions; the global diagram therefore differs qualitatively from the RN scenario.
Further details and the Penrose diagram can be found in Ref. [87].
We note that $q$ plays a central role in the EMP theory, functioning not only as a regularization parameter for constructing wormholes and resolving spacetime singularity, but also as a geometric parameter determining the radius of the wormhole throat. It is further interpreted as a scalar charge, in analogy to electric charge. An interesting question: Is there a physically motivated range of $q$, for instance, from the energy conditions? The answer is negative. The phantom scalar field that sources the EMP solutions explicitly violates the null energy condition (NEC)  [82, 83], and such a condition is a prerequisite for constructing traversable wormholes. Moreover, the NEC is the weakest of the four energy conditions. That is, if the NEC is violated, the other three energy conditions, the weak, strong, and dominant energy conditions, are also violated. Therefore, the energy conditions cannot be used to impose meaningful constraints on the scalar charge $q$. On the other hand, the EMP framework, as a result of the four-dimensional low-energy effective theory of string theory, must recover general relativity in the classical limit. This is achieved under the parameter choice,
$\gamma_{1}=\frac{1}{4},\qquad\gamma_{2}=\frac{8qM}{Q^{2}},\qquad q\to 0^{+},$ (15)
which implies that the quantum-gravity consistency sets a lower bound for $q$, namely zero. This limit is consistent with fundamental principles of general relativity, e.g., the preservation of U(1) gauge symmetry, enforcement of the no-scalar-hair theorem, and satisfaction of the positive mass theorem [95], thereby establishing theoretical consistency of the EMP theory with the Einstein-Maxwell dynamics.
@@ 3.1 General set up
The charged EB wormhole solution of the EMP theory [87], as a special case of Eq. (19), has the following coefficients,
@@
$e_{0\mu}$ ${}={}$ $\bigl($ $\ \ -\mathrm{e}^{\nu},$ $0,$ $\quad\qquad 0,$ $\quad 0\bigr),$
$e_{1\mu}$ ${}={}$ $\bigl($ $-\sigma\mathrm{e}^{\psi},$ $\mathrm{e}^{\psi},$ $\ -q_{2}\mathrm{e}^{\psi},$ $\ -q_{3}\mathrm{e}^{\psi}\bigr),$
$e_{2\mu}$ ${}={}$ $\bigl($ $\ \qquad 0,$ $\ \qquad 0,$ $\,\qquad\mathrm{e}^{\mu_{2}},$ $\quad 0\bigr),$
$e_{3\mu}$ ${}={}$ $\bigl($ $\ \qquad 0,$ $0,$ $\quad\qquad 0,$ $\quad\mathrm{e}^{\mu_{3}}\bigr).$ (22)
@@ 3.2 Gravito-Electromagnetic field perturbation
For gravito-electromagnetic field perturbations, the linearization of the Einstein equation Eq. (7) yields
@@
The left-hand side of the above equation can be computed via Cartan’s equations of structure applied to the 1-forms in Eq. (22). In particular, the axial components $\delta R_{12}$ and $\delta R_{13}$ take the form,
@@
where $\mathcal{Q}(r,\theta,t)=\varrho^{2}\mathrm{e}^{2\nu}Q_{23}\sin^{3}\theta$ encodes gravitational field perturbations and $B(r,\theta)=\delta P_{01}\sin{\theta}$ represents electromagnetic field perturbations. Note that those terms involving $\delta\left(\partial_{a}\phi\partial_{b}\phi\right)$ vanish in Eqs. (29) and (30) by virtue of the axial symmetry of the background metric Eq. (19). Additionally, the phantom scalar (spin-0) perturbations are parity-even and therefore appear only in the polar (even-parity) sector in the static and spherically symmetric EMP spacetime; they do not source the axial (odd-parity) equations.
Returning to the axial sector and assuming a time dependence $\mathrm{e}^{-\mathrm{i}\omega t}$ in the perturbations, elimination of $\sigma$ in Eqs. (29) and (30) yields the first master equation for gravito-electromagnetic field perturbations,
@@
$\left[Z^{-1}\mathrm{e}^{2\nu}\left(Z\varrho\mathrm{e}^{\nu}B\right)_{,2}\right]_{,2}+\frac{\mathrm{e}^{\nu}}{\varrho}\left(\frac{B_{,3}}{\sin{\theta}}\right)_{,3}\sin{\theta}+\left(\omega^{2}\varrho\mathrm{e}^{-\nu}-\frac{Q^{2}}{\varrho^{3}}Z\mathrm{e}^{\nu}\right)B=-Q\frac{\mathcal{Q}_{,3}}{\varrho^{4}\sin{\theta}}.$ (33)
@@
so that $H_{1}$ and $H_{2}$ represent electromagnetic and gravitational field perturbations, respectively, with subscripts denoting the spin of the fields. We further adopt the tortoise coordinate $r_{*}$ defined by
@@
$\mathbf{H}=\left(\begin{array}[]{c}H_{1}\\ H_{2}\end{array}\right),\qquad\mathbf{V}=\left(\begin{array}[]{cc}V_{11}&V_{12}\\ V_{21}&V_{22}\end{array}\right),$ (47)
@@
In the RN background, the effective potential matrix in the master equation can be diagonalized by a constant matrix, yielding decoupled perturbation equations. However, the coupling cannot be removed in this manner [^1] It has come to our attention that some studies [97, 98, 99] have inappropriately applied the RN decoupling procedure in their work. To avoid further misleading, we explicitly clarify here that the RN decoupling procedure is not valid in the EMP spacetime. in the EMP spacetime, so we shall apply numerical methods directly to the coupled system.
While the primary focus of this work is on the axial perturbation of the EMP spacetime, a brief discussion of the polar sector is warranted to provide a broader context for the stability analysis. A complete characterization of the linear dynamics requires the investigation of both sectors. Below we present a concise qualitative discussion of polar perturbations. We also include a approximate quantitative treatment to provide a preliminary outlook on the polar sector.
3.3 Phantom scalar field perturbation
In a spherically symmetric background, perturbations can be decomposed to the sectors with a definite parity. The axial sector (odd-parity), which we have studied in detail, is associated with frame dragging and with coupled gravito-electromagnetic degrees of freedom. By contrast, the polar sector (even-parity) contains a richer set of dynamics, including radial pulsations of the spacetime geometry, variations in the matter distribution (energy–momentum tensor), and the even-parity components of the Maxwell field.
Within the EMP theory, the polar perturbation is expected to be considerably more complex than its axial counterpart. This increased complexity stems from the fact that all background fields — the EMP metric field $g_{\mu\nu}$, the Maxwell field $A_{\mu}$, and the phantom scalar field $\phi$ — are of polar type; accordingly, their linear perturbations will inherently couple to one another, leading to a gravito-electromagnetic-phantom scalar coupled system with dynamics distinct from that of the axial sector. In particular:
  • In addition to the gravito-electromagnetic degrees of freedom, the polar spectrum contains extra modes sourced by perturbations of the phantom scalar field. The phantom scalar field perturbation is influenced only by the perturbed electromagnetic field. By contrast, the electromagnetic field perturbations are influenced by both the gravitational and phantom scalar field perturbations, and similarly, the gravitational perturbations are influenced by both the electromagnetic and phantom scalar field perturbations, leading to a fully coupled system.
  • The effective potentials in the polar sector may have qualitatively different structures, which can change stability thresholds for wormholes and shift quasinormal frequencies relative to the axial sector.
  • The breaking of isospectrality between the axial and polar modes has been reported in other theories with scalar hairs (see, e.g., Ref. [100, 101]), because the scalar field enters the polar equations but not the axial ones; similar isospectral-breaking is therefore foreseeable in the EMP spacetime.
A comprehensive analysis of polar perturbations in the EMP spacetime is technically challenging and is left for future work. Here we adopt a controlled approximation in which linear electromagnetic perturbations are no longer treated as independent dynamical degrees of freedom in the polar sector, but are assumed to arise predominantly from the implicit dependence of the background field on the geometric parameter $q$.
By linearizing the equation of motion Eq. (5), we obtain the master equation of phantom scalar field perturbations,
$\square\delta\phi$ $=-\frac{1}{4}\left(\frac{\partial^{2}Z^{-1}}{{\partial\phi}^{2}}F^{2}\delta\phi+\frac{\partial Z^{-1}}{\partial\phi}\delta F^{2}\right).$ (49)
The right-hand side of the above equation consists of the linear term of the phantom scalar field $\delta\phi$ and that of the electromagnetic invariant $\delta F^{2}$, revealing a coupling between the perturbed phantom scalar field and the perturbed Maxwell field. For the EMP background, the first electromagnetic invariant reads
$F^{2}=F^{\mu\nu}F_{\mu\nu}=-2F_{rt}^{2}=-2\left(\frac{QZ}{r^{2}+q^{2}}\right)^{2}.$ (50)
Analogous to the metric components in Eq. (21), the linear perturbation of the Maxwell field in the EMP spacetime can be ascribed to an increment of $F_{rt}$ together with the non-vanishing of the other components of the electromagnetic tensor. Importantly, the variation of the electromagnetic invariant is governed solely by $\delta F_{rt}$ at the linear order,
$\delta F^{2}=-4F_{rt}\delta F_{rt}=-4F_{rt}\frac{\partial F_{rt}}{\partial q}\frac{\partial q}{\partial\phi}\delta\phi,$ (51)
where the contributions from the other components enter only at higher orders. In the final equality we adopt the approximation that the variation $\delta F_{rt}$ is dominated by the implicit dependence of $F_{rt}$ on the geometric parameter $q$, and we hold the total electric charge fixed ($\delta Q=0$), consistent with source-free perturbations that preserve the asymptotic condition. Under this assumption only the $q$-dependence is retained when converting $\delta F_{rt}$ into $\delta\phi$ via the chain rule, which provides an effective closure of the phantom scalar-electromagnetic coupling at the linear order. This closure is an approximation and should be relaxed when polar Maxwell degrees of freedom or charged sources are dynamically important.
with the effective potential
$V=h\left(\frac{rh^{\prime}}{\varrho^{2}}+\frac{q^{2}h}{\varrho^{4}}+\frac{l(l+1)}{\varrho^{2}}-\frac{Q^{2}}{8q^{2}\varrho^{2}}\left(2rZ^{\prime}-\varrho^{2}Z^{\prime\prime}\right)\right),$ (54)
4 Instabilities of wormholes